OpenStudy (anonymous):

The problem: log(x^(3)y^(2))-2logx√3:(y)-3log((x)/(y)) My answer: log((y^((14)/(3)))/(x^(2))) Book's answer: log((y^((13)/(3)))/(x^(2))) What am I doing wrong?

OpenStudy (anonymous):

i am a bit confused as to the original problem it is \[\log(x^2y^2)-2\log(x\sqrt[3]{y})-3\log(\frac{x}{y})\]?

OpenStudy (anonymous):

Almost. The first x^{3}

OpenStudy (anonymous):

is not x^{2}

OpenStudy (anonymous):

\[\log(x^3y^2)-2\log(x\sqrt[3]{y})-3\log(\frac{x}{y})\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and you want to combine in to a single log right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok then it is mostly bookkeeping with the exponents

OpenStudy (anonymous):

\[\log(x^3y^2)-2\log(x\sqrt[3]{y})-3\log(\frac{x}{y})\] \[\log(x^3y^2)-\log(x^2\sqrt[3]{y^2})-\log(\frac{x^3}{y^3})\]

OpenStudy (anonymous):

the input is \[\frac{x^3y^2y^3}{x^2y^{\frac{2}{3}}x^3}\]

OpenStudy (anonymous):

simplifies to \[\frac{y^{\frac{13}{3}}}{x^2}\]

OpenStudy (anonymous):

oh that is what you got! i think your answer is right

OpenStudy (anonymous):

Ok now I see where my mistake was. I didn't distribute the 2 in front of the log, only applied it to the x.

OpenStudy (anonymous):

I was off and I couldn't figure it out why. Now i know.

OpenStudy (anonymous):

oh that is the book answer ok i hope it is clear

OpenStudy (anonymous):

yes, thank you very much for your help

OpenStudy (anonymous):

yw